# Hyperbolic Random Graphs: Clique Number and Degeneracy with Implications for Colouring

### v1

@@ Appendix B Bounds on $\theta_{R}(\cdot,\cdot)$
$\mathbb{E}\left[X_{v}\right]\geq\mathbb{E}\left[|\operatorname{\Gamma}(u^{*})|\right]\in n^{\Omega(1)}.$ (by (G.1) and (14))
@@
See 5.3
Proof G.2.
To calculate the probability that a vertex lies in the inner-neighbourhood of another vertex we notice that, independent of the geometric distance, a vertex with weight $w$ is adjacent to any vertex with weight $w^{\prime}$ if $w^{\prime}\geq\frac{n}{w\lambda}$ since $t(w,w^{\prime})=\frac{\lambda ww^{\prime}}{2n}\geq\frac{\lambda w\frac{n}{w\lambda}}{2n}=\frac{1}{2}$ and the maximal distance between two points in the torus $\mathbb{T}^{1}=[0,1]$ is at most $1/2$ Thus, using $\beta=2\alpha+1$, the probability that a vertex $v$ is in the inner-neighbourhood of a vertex $u$ with weight $w$ is given by
$\mathbb{P}\left(v\in\Gamma(u)\right)$ $=\mathbb{P}\left(u\sim v\cap W\geq w|w_{u}=w\right)$
[…]
This concludes the first step of the proof. Next, we are interested in the value $w^{*}$, which maximizes the expected inner-degree. To this end, we consider the probability measure of the inner-neighbourhood, take its derivative with respect to $w$ and set it equal to 00. This reveals
Differentiating yields
$\frac{\mathrm{d}}{\mathrm{d}w}\mathbb{P}\left(v\in\Gamma(u)\right)$ $=\frac{2(nw)^{-(2\alpha+1)}(1-\alpha)\alpha\left(n^{2\alpha}w^{2}\lambda-nw^{4\alpha}\lambda^{2\alpha}\right)}{2\alpha-1}$
Solving for the maximum yields
$w^{*}=(2(1-\alpha))^{\frac{1}{4\alpha-2}}\sqrt{\frac{n}{\lambda}}.$ (16)
We plug in $w^{*}$ for the weight of $u$ denoted by $u^{*}$ into $\mathbb{P}\left(v\cap\Gamma(u^{*})\right)$ and get by (G.2) that
Recalling that $\sigma(G)=(1\pm o(1))\lambda^{-\alpha}n^{1-\alpha}$ w.e.h.p., the upper bound now follows from the expectation of $|\operatorname{\Gamma}(u^{*})|$ and applying a Chernoff bound in conjunction with a union bound. The lower bound is established by showing that there exists a vertex within the range of weights $\tilde{w}\in[w^{*}(1+n^{\alpha-1}\log^{2}(n))^{-1/(2\alpha)},w^{*}]$ w.e.h.p. Using the Pareto distribution and $w^{*}\in\Theta(\sqrt{n})$, we calculate the probability for a vertex $u$ to belong to the range of weights $\tilde{w}$
$\mathbb{P}\left(w_{u}\in\tilde{w}\right)$ $=\mathbb{P}\left(w^{*}(1+n^{\alpha-1}\log^{2}(n))^{-1/(2\alpha)}\leq W_{u}\leq w^{*}\right)$
$=1-(w^{*})^{-2\alpha}-(1-(w^{*}(1+n^{\alpha-1}\log^{2}(n))^{-1/(2\alpha)})^{-2\alpha})$
$=(w^{*})^{-2\alpha}n^{\alpha-1}\log^{2}(n))$
$=\Theta(1)\frac{\log^{2}(n)}{n}.$
By this we have $\mathbb{E}\left[|V\cap\tilde{w}|\right]\in\omega(\log(n))$. Using a Chernoff bound there exists a vertex within the weight range $\tilde{w}$ w.e.h.p. To conclude the proof we lower bound the inner-degree a vertex included in the weight range $\tilde{w}$ and denote it by $\tilde{u}$. Note that $\tilde{w}=(1-o(1))w^{*}=(2-o(1)(1-\alpha))^{\frac{1}{4\alpha-2}}\sqrt{\frac{n}{\lambda}}$. We then have via G.2

### v2

@@ A.5 Geometric inhomogeneous random graphs
$\mathbb{E}\left[X_{v}\right]\geq\mathbb{E}\left[|\operatorname{\Gamma}(u^{*})|\right]\in n^{\Omega(1)}.$ (by (A.12) and (15))
@@
See 5.5
Proof A.13.
Fix any $\alpha\in(1/2,1)$ as given in the statement. Draw the hyperbolic random graph $H\sim\mathcal{G}(n,\alpha,C)$ and the geometric inhomogeneous random graph $G\sim\mathcal{G}(n,2\alpha+1,\lambda)$. Using Theorem 5.3 for the degeneracy of the GIRG we obtain w.e.h.p.
$\kappa(G)/\sigma(G)\geq(2-o(1))(2(1-\alpha))^{({1-\alpha})/{(2\alpha-1)}}.$
In contrast, upper bounding the degeneracy of the HRG with Theorem 3.7 reveals w.e.h.p.
$\kappa(H)/\sigma(H)\leq((4/3)^{\alpha}-o(1)).$
Combining the two inequalities, it now follows that $\kappa(G)/\sigma(G)-\kappa(H)/\sigma(H)$ is bounded away from 00 w.e.h.p., which is what we sought out to show.